on January 1, 2018, Doty Co. redeemed its 10-year bonds of $4,000,000 par value for 103....

  1. on January 1, 2018, Doty Co. redeemed its 10-year bonds of $4,000,000 par value for 103. They were originally issued on January 1, 2006 at 95 with a maturity date of
    January 1, 2016. Doty amortizes discounts and premiums using the straight-line method.
    1. What amount of loss should Doty recognize on the redemption of these bonds (ignore taxes)?
Answer #1

Redemption value = 4000000*1.03 = 4120000

Bonds payable = 4000000

Cash = 4000000*0.95 = 3800000

Discount on bonds payable = 4000000-3800000 = 200000

Amount of loss Doty should recognize on the redemption of bonds = 4120000-(3800000+200000) = $120000